# latexYou have requested package

### v1

@@ 4 Results
For all experiments in this paper, we used the YOLOv5n model [11]. Initial benchmarking was conducted on the Pascal VOC dataset [9] to avoid training multiple models on the larger COCO dataset [14], with final benchmarking results generated using COCO. Tables 1 and 2 summarize mAP, RAM (NPU only), and latency values for both the reference models and our mixed activation models. The baseline models consist of YOLO5 trained on COCO using SiLU, ReLU, and Hardswish activations, respectively. The LZCM1 model aims to minimize overall latency by mixing SiLU and ReLU activations using a simple search approach, while the LZCM2 model prioritizes maximizing accuracy by combining the more accurate activations, SiLU and Hardswish. Both models achieve mAP values slightly lower, but very close to, those of the SiLU and Hardswish reference models. In terms of latency, the LZCM1 model performs slightly better than the SiLU model across all devices (e.g., 21.87 ms vs. 22.35 ms on the NPU). The SiLU/Hardswish mixed activation model demonstrates a better accuracy-latency trade-off, being 12.5% faster than the SiLU model but 5.3% slower than the Hardswish model on the NPU. A similar trend is observed on A53 and A57 CPUs as well as the Jetson Nano GPU. Lastly, the naive approach of replacing the first three activations from SiLU to ReLU resulted in slightly slower performance compared to other baseline models, but the drop in mAP was minimal (0.002) compared to the SiLU model.
@@
Figure 4: Performance of reference models and activation NAS generated models on different devices.
@@
ActNAS1 and ActNAS3 were generated using our Activation NAS approach, targeting the NPU as the primary device, as discussed in the previous section. ActNAS1 achieved the highest mAP and the lowest latency, showing 6.26%, 3.32%, 0.97%, and 1.69% improvements in efficiency on the NPU, Jetson Nano GPU, Cortex A57 CPU, and A53 CPU, respectively, compared to the Hardswish model. It also outperformed the SiLU models by 22.28%, 8.90%, 11.92%, and 13.64% on the same devices. Additionally, we found that ActNAS1 required 58% less memory than the SiLU model on the NPU. These experiments demonstrate that the performance of YOLO models can be preserved by creating mixed activation models using a zero-cost proxy-based accuracy estimation, along with latency and memory computation on the target device, paired with an effective NAS approach. ActNAS2, generated with the Jetson Nano GPU as the target, outperformed both the baseline and Hardswish models on Jetson Nano, highlighting the hardware-aware nature of our search method, as it adapts to different hardware profiles. We also observed that the best latency-accuracy trade-off models tended to use ReLU variants (ReLU6, LeakyReLU, and ReLU) in the initial layers. None of the ActNAS models utilized SiLU or Hardswish in the first three layers. This suggests that replacing the initial layers with more efficient activations and following them with high-performance activations—selected via NAS to minimize accuracy loss—produces optimal models across a range of hardware devices. Figure 4 compares our models with baseline models, showing that Activation NAS models offer the best accuracy-latency trade-off on all tested hardware.
@@ 5 Conclusion
In this paper, we introduced a novel approach called Activation NAS, which generates YOLO model architectures using a search space of different activations for individual constituent layers. Our mixed activation models demonstrate slightly better performance than the SiLU-based YOLO model on the COCO dataset, while being significantly more efficient, with 6% to 23% lower latency depending on the hardware. The NAS approach leverages the zero-cost NWOT score to avoid training a large set of candidate models, making this method practical for model optimization. These experiments suggest that, instead of using a single activation function across the entire model, a combination of latency and accuracy budgets can be used to create custom mixed activation models that fit within specified accuracy, latency, and memory constraints. Future work will expand the benchmarking results to YOLO models with greater width and depth, allowing us to further explore performance improvements and trends in speedup.

### v2

@@ 4 Results
We conducted our experiments on YOLO5n and YOLO8m [8] models and the selection is to cover two family of YOLO models with different levels of complexity. Initial benchmarks were performed on YOLO5n model for Pascal VOC dataset [5] to avoid training multiple models on the larger COCO dataset [12]. Still, all the final benchmarking is performed on COCO dataset. Tables 1 and 2 summarize mAP, RAM (NPU only), and latency values for both the reference models and our mixed activation models for YOLO5n model. The baseline models consist of COCO-trained YOLO5 using SiLU, ReLU, and Hardswish activations, respectively. The LZCM1 model aims to minimize overall latency by mixing SiLU and ReLU activations using a simple search approach, while the LZCM2 model prioritizes maximizing accuracy by combining the more accurate activations, SiLU and Hardswish. Both models achieve mAP values slightly lower, but very close to SiLU and Hardswish reference models. In terms of latency, the LZCM1 model performs slightly better than the SiLU model across all devices (e.g., 21.87 ms vs. 22.35 ms on the NPU). The SiLU/Hardswish mixed activation model demonstrates a better accuracy-latency trade-off, being 12.5% faster than the SiLU model but 5.3% slower than the Hardswish model on the NPU. A similar trend is observed on A53 and A57 CPUs, as well as the Jetson Nano GPU. Lastly, the naive approach of replacing the first three activations from SiLU to ReLU resulted in slightly slower performance compared to other baseline models, but the drop in mAP was less than 1% compared to the SiLU model.
@@
Figure 6: Performance of YOLO5n reference models and ActNAS models on different edge devices.
Figure 7: Performance of YOLO8m reference models and ActNAS models on different edge NPUs.
@@
ActNAS5n.1 and ActNAS5n.2 were generated using our activation NAS approach, targeting the NPU as the primary device, as discussed in the previous section. The mAP of ActNAS5n.1 model is within 1% drop compared to both SiLU and HardSwish variants but this model has lowest latency showing 6.26%, 3.32%, 0.97% and 1.69% improvements in efficiency on the NPU1, Jetson Nano GPU, Cortex A57 CPU and A53 CPU, respectively, compared to the Hardswish model. It also outperformed the SiLU model by 22.28%, 8.90%, 11.92%, and 13.64% on the these devices. Furthermore, we found that ActNAS5n.1 required 58% less memory than the SiLU model on NPU1. ActNAS5n.2, generated with the Jetson Nano GPU as the target, outperformed both the baseline and HardSwish models on Jetson Nano, highlighting the hardware-aware nature of our search method, as it adapts to different hardware profiles. We also observed that the best latency-accuracy trade-off models tends to use ReLU variants (ReLU6, LeakyReLU, and ReLU) for initial layers. None of the ActNAS models utilized SiLU or Hardswish in the first two layers. This suggests that replacing the initial layers with more efficient activations and following them with high-performance activations selected via NAS to minimize accuracy loss produces optimal models across a range of hardware devices. Figure 6 compares our models with baseline models, showing that Activation NAS models offer the best accuracy-latency trade-off on all tested hardware. In order to test the scalability of this approach, we repeated these experiments with YOLO8m models which has more activation layers compared to YOLO5n. We observerd that the impact of different activations is more prominent on NPUs as latency is not only affected by the complexity of operations but also by the memory usage. Therefore, we focused all the YOLO8m experiments on three different NPUs. Table 3 shows the mAP/latency trade-off of ActNAS8m.1 and ActNAS8m.2 models compared to baseline models. These models are created to be NPU2 hardware aware and it is evident that ActNAS8m.2 model is 1.66x faster compared to the SiLU model without significant impact on accuracy of the model. Figure 7 shows the mAP-latency curve for ActNAS YOLO8m models compared to baseline model.
5 Hardware Awareness
As mentioned in section 3, all of the ActNAS models use one of the edge device for latency/memory computation. Since the latency and memory impact of each change in the model correlates to the hardware architecture, memory limitations and compiler optimization, we expect the resulting optmized model to have the best performance on that hardware. In other words, the model is ”device” aware. For example, ActNAS5n.1, ActNAS5n.2, ActNAS8m.1 and ActNAS8m.2 are NPU1, Jetson Nano GPU, NPU2 and NPU2 aware models. The bar chart in Figure 8 shows the speedup of NPU1 aware ActNAS5n.1 and NPU2 aware ActNAS8m.2 speed up on five different devices. NPU1 and NPU2 have similar architecture but with different memory configuration. We can see that both NPU aware models provide highest speedup on these devices but very low speedup on other devices. These results also suggest that the NAS is able to learn model configuration suited best for these devices.
Figure 8: ActNAS model speedup on different hardware where the model is optimized for NPU1 (NPU1 and NPU2 have similar configuration)
@@ 6 Conclusion
In this paper, we present a novel approach called Activation NAS (ActNAS), designed to generate YOLO model architectures through a unique search space of varied activation functions for each layer. Our mixed-activation models deliver performance comparable to SiLU/Hardswish-based models on the COCO dataset, while offering up to 1.66x greater efficiency, depending on the hardware. This efficiency is achieved using a Neural Architecture Search (NAS) approach that leverages the zero-cost NWOT score, enabling us to skip the costly process of training a vast number of candidate models, thus making model optimization more practical. Our experiments suggest that, rather than using a single activation function across all model layers, a combination of latency and accuracy constraints can guide the creation of custom mixed-activation models tailored to meet specific accuracy, latency, and memory requirements. Furthermore, we demonstrate that models optimized with particular hardware in mind (hardware-aware models) achieve peak performance on that hardware, yet may underperform on others. This finding is crucial for model accelerators and NPUs with varying architectures, where hardware-agnostic optimizations often fail to yield efficient models.